# Benchmarking Retrieval-Augmented Multimomal Generation for Document Question Answering

### v1

@@ Abstract
Document Visual Question Answering (DocVQA) faces dual challenges in processing lengthy multimodal documents (text, images, tables) and performing cross-modal reasoning. Current document retrieval-augmented generation (DocRAG) methods remain limited by their text-centric approaches, frequently missing critical visual information. The field also lacks robust benchmarks for assessing multimodal evidence selection and integration. We introduce MMDocRAG, a comprehensive benchmark featuring 4,055 expert-annotated QA pairs with multi-page, cross-modal evidence chains. Our framework introduces innovative metrics for evaluating multimodal quote selection and enables answers that interleave text with relevant visual elements. Through large-scale experiments with 60 VLM/LLM models and 14 retrieval systems, we identify persistent challenges in multimodal evidence retrieval, selection, and integration. Key findings reveal advanced proprietary LVMs show superior performance than open-sourced alternatives. Also, they show moderate advantages using multimodal inputs over text-only inputs, while open-source alternatives show significant performance degradation. Notably, fine-tuned LLMs achieve substantial improvements when using detailed image descriptions. MMDocRAG establishes a rigorous testing ground and provides actionable insights for developing more robust multimodal DocVQA systems. Our benchmark and code are available at https://mmdocrag.github.io/MMDocRAG/.
@@ 1 Introduction
DocVQA [43] focuses on visual question answering over documents with rich multimodal content. Multimodal documents (e.g., financial reports, technical manuals, and medical records) present significant challenges for DocVQA: (i) they are typically lengthy, complicating the identification of key evidence, and (ii) they require complex reasoning across various modalities, including images, tables, charts, and layout structures. Thus, recent studies [10, 61, 9] have adopted document retrieval-augmented generation (DocRAG), which first retrieves relevant document pages and then generates answers by selecting and composing supporting evidence. However, current DocRAG systems show significant limitations, resulting in perspective narrowing, as highlighted in Table 1: 1. Unimodal Bias: Generated answers frequently over-rely on plain text, neglecting valuable visual information such as charts and tables. Prior work [83, 42] has shown that multimodal content greatly enhances user understanding, supporting the notion that “a single image is worth a thousand words”. Moreover, presenting multimodal evidence improves answer traceability and credibility by enabling users to directly verify supporting information. 2. Evaluation Flaws: Existing benchmarks [41, 9, 16] primarily assess the recall of retrieved quotes or the quality of textual answers. There are no benchmarks for evaluating a model to (i) select relevant multimodal evidence from noisy retrieved quotes or (ii) align and integrate multimodal content with text in a coherent and logical manner. These gaps hinder the evaluation in complex multimodal RAG scenarios.
@@
In response to these challenges, we propose MMDocRAG, a comprehensive multimodal document question answering benchmark, with an annotation exemplified in Figure LABEL:fig:teaser. MMDocRAG consists of 4,055 expert-annotated question-answer pairs, each accompanied by multimodal evidence chains which may span multiple pages and modalities, including both text and image quotes. Evidence is provided at multiple granularities, ranging from coarse-grained page-level screenshots to fine-grained quotes extracted based on document layout. In addition to these annotations, MMDocRAG introduces two novel evaluation features: (1) Quote Selection: We propose a practical evaluation metric that measures a model’s ability to select and integrate relevant multimodal quotes. To increase task difficulty, we include hard text and image negatives  [^1] Hard negatives refer to quotes retrieved with high textual or visual similarity but irrelevant to the question. mixed with gold (relevant) quotes. (2) Multimodal Output Paradigm: Our benchmark supports multimodal answers, allowing document figures, infographics, charts, and tables to be interleaved within textual responses. This paradigm enhances both the interpretability and cognitive effectiveness of generated answers.
@@
Utilizing MMDocRAG, we conduct comprehensive experiments on DocVQA/RAG tasks. Our study includes 60 latest large models, among which 33 VLMs can handle multimodal (interleaved text and image) inputs and 27 LLMs can only process text inputs. For multimodal tasks with LLMs, we either extract text from images using OCR [65] tools (“OCR-text”) or use VLMs [49, 54] to generate detailed image descriptions (“VLM-text”). We fix the number of input quotes to 15 or 20 for multimodal generation. Experimental results highlight the complexities of multimodal DocRAG: the best model, GPT4.1 [52], achieves an F1 score of only 70.2% for quote selection. For multimodal answer quality, we assess fluency, citation quality, text-image coherence, reasoning, and factual accuracy, with GPT4.1 achieving the highest scores. Overall, proprietary VLMs significantly outperform open-sourced VLMs and LLMs. The advanced proprietary VLMs generally show better performance using multimodal inputs over pure-text inputs, and the performance gap is modest. In contrast, open-source or smaller proprietary VLMs show significant performance boost using pure-text inputs than multimodal inputs. Meanwhile, fine-tuning Qwen2.5-instruct LLMs [59] (3–72B parameters) yields substantial performance improvements. Notably, LLMs leveraging VLM-text significantly outperform those using OCR-text. Additionally, we evaluate the retrieval performance of 6 text, 4 visual, and 4 hybrid retrievers, further highlighting the challenges of extracting relevant multimodal quotes from long documents. In summary, our contributions are:
@@
  • We propose MMDocRAG benchmark for evaluating multimodal generation on DocVQA/RAG tasks. Our dataset include over 4,000 QA pairs, diverse forms of evidence, a mixture of gold and noisy quotes to enable nuanced quote selection, and answers with interleaved multimodal content.
@@
  • We conduct extensive evaluations on multimodal RAG, covering (i) retrieval performance on 6 text, 4 visual,4 hybrid retrievers, (ii) quote selection F1 and (iii) multimodal answer quality across 37 open-source and 23 proprietary models, and 5 models finetuned using MMDocRAG dev-set.
@@ 2.1 Construction
We utilize the document corpus from the MMDocIR dataset [16], which consists of 313 documents spanning over 10 diverse domains. These documents are sufficiently long (averaging 65.1 pages) and provide rich multimodal coverage. We process the documents with MinerU [71], which leverages LayoutLMv3 [27] to detect page layouts and classify them as body text, titles, equations, figures, tables, etc. Each identified layout serves as a content-aware chunk, or “quote”. Text quotes correspond to layouts such as equations or paragraphs, and are stored in text format. For image quotes (e.g., tables or figures), we extract text using OCR [65] (“OCR-text”) and generate detailed descriptions using VLMs [49, 54] (“VLM-text”). Consequently, each image quote is stored in three formats: original image, OCR-text, and VLM-text. After indexing all documents, we carefully select pages with rich multimodal and text information. This process yields 2,373 high-quality pages, forming the basis for subsequent annotation.
@@ 3 Task Definition
Let $\mathcal{D}$ denote a document corpus consisting of text quotes $\mathcal{T}=\{t_{1},t_{2},\dots,t_{m}\}$ and image quotes =$\mathcal{I}=\{i_{1},i_{2},\dots,i_{n}\}$, as extracted via layout detection (see Section 2.1). On average, documents in MMDocRAG contain 63 image quotes and 536 text quotes. The objective is to retrieve a subset of quotes that are most relevant to a query $Q$ from $\mathcal{T}$ and $\mathcal{I}$, by ranking them based on similarity scores, $\textrm{Sim}(Q,t)$ and $\textrm{Sim}(Q,i)$. The top-$k$ quotes, where $k\ll n+m$, are selected as candidate evidence.

### v2

# Benchmarking Retrieval-Augmented Multimodal Generation for Document Question Answering

@@ Abstract
Document Visual Question Answering (DocVQA) faces dual challenges in processing lengthy multimodal documents (text, images, tables) and performing cross-modal reasoning. Current document retrieval-augmented generation (DocRAG) methods remain limited by their text-centric approaches, frequently missing critical visual information. The field also lacks robust benchmarks for assessing multimodal evidence selection and integration. We introduce MMDocRAG, a comprehensive benchmark featuring 4,055 expert-annotated QA pairs with multi-page, cross-modal evidence chains. Our framework introduces innovative metrics for evaluating multimodal quote selection and enables answers that interleave text with relevant visual elements. Through large-scale experiments with 60 VLM/LLM models and 14 retrieval systems, we identify persistent challenges in multimodal evidence retrieval, selection, and integration. Key findings reveal advanced proprietary LVMs show superior performance than open-sourced alternatives. Also, they show moderate advantages using multimodal inputs over text-only inputs, while open-source alternatives show significant performance degradation. Notably, fine-tuned VLM/LLMs achieve substantial improvements for multimodal generation. MMDocRAG establishes a rigorous testing ground and provides actionable insights for developing more robust multimodal DocVQA systems. Our benchmark and code are available at https://mmdocrag.github.io/MMDocRAG/.
@@ 1 Introduction
DocVQA [49] focuses on visual question answering over documents with rich multimodal content. Multimodal documents (e.g., financial reports, technical manuals, and medical records) present significant challenges for DocVQA: (i) they are typically lengthy [18], complicating the identification of key evidence, and (ii) they require complex reasoning across various modalities, including images, tables, charts, and layout structures. Thus, recent studies [10, 67, 9, 30, 20] have adopted document retrieval-augmented generation (DocRAG), which first retrieves relevant document pages and then generates answers by selecting and composing supporting evidence. However, current DocRAG systems show significant limitations, resulting in perspective narrowing, as highlighted in Table 1: 1. Unimodal Bias: Generated answers frequently over-rely on plain text, neglecting valuable visual information such as charts and tables. Prior work [91, 48] has shown that multimodal content greatly enhances user understanding, supporting the notion that “a single image is worth a thousand words”. However, current AIGC models struggle in generating informative visualizations from scratch. Ideally, the multimodal evidence displayed in answer must be traceable and credible, allowing users to verify supporting information as shown in Figure 1. 2. Evaluation Flaws: Existing benchmarks [47, 9, 19] primarily assess the recall of retrieved quotes or the quality of textual answers. There are no benchmarks for evaluating a model to (i) select relevant multimodal evidence from noisy retrieved quotes or (ii) align and integrate multimodal content with text in a coherent and logical manner. These gaps hinder the evaluation in complex multimodal RAG scenarios.
Figure 1: Comparison of different QA paradigms: text-only, image-only, and interleaved text-image.
@@
In response to these challenges, we propose MMDocRAG, a comprehensive multimodal document question answering benchmark ($\S$2), with an annotation exemplified in Figure LABEL:fig:teaser. MMDocRAG consists of 4,055 expert-annotated question-answer pairs, each accompanied by multimodal evidence chains which may span multiple pages and modalities, including both text and image quotes. Evidence is provided at multiple granularities, ranging from coarse-grained page-level screenshots to fine-grained quotes extracted based on document layout. In addition to these annotations, MMDocRAG introduces two novel evaluation features: (1) Quote Selection: We propose a practical evaluation metric that measures a model’s ability to select and integrate relevant multimodal quotes. To increase task difficulty, we include hard text and image negatives [^1] Hard negatives refer to quotes retrieved with high textual or visual similarity but irrelevant to the question. mixed with gold (relevant) quotes. (2) Multimodal Output Paradigm: Our benchmark supports multimodal answers, allowing document figures, infographics, charts, and tables to be interleaved within textual responses. This paradigm enhances both the interpretability and cognitive effectiveness of generated answers.
@@
Utilizing MMDocRAG, we conduct comprehensive experiments on DocVQA/RAG tasks. Our study includes 60 latest large models, among which 33 VLMs can handle multimodal (interleaved text and image) inputs and 27 LLMs can only process text inputs. For multimodal tasks with LLMs, we either extract text from images using OCR [71] tools (“OCR-text”) or use VLMs [55, 60] to generate detailed image descriptions (“VLM-text”). We fix the number of input quotes to 15 or 20 for multimodal generation. Experimental results ($\S$4.3) highlight the complexities of multimodal DocRAG: the best model, GPT4.1 [58], achieves an F1 score of only 70.2% for quote selection. For multimodal answer quality, we assess fluency, citation quality, text-image coherence, reasoning, and factual accuracy, with GPT4.1 achieving the highest scores. Overall, proprietary VLMs significantly outperform open-sourced VLMs and LLMs. Meanwhile, fine-tuning Qwen2.5-instruct LLMs [65] (3–72B parameters), Qwen2.5-VL-Instruct VLMs [3] (3&7B), and InternVL-3 VLMs [90] (8&9B) yields substantial performance improvements. It is worthnoting that the advanced proprietary VLMs generally show better performance using multimodal inputs over pure-text inputs, and the performance gap is modest. In contrast, open-source or smaller proprietary VLMs show significant performance boost using pure-text inputs than multimodal inputs ($\S$4.4). Notably, LLMs leveraging VLM-text significantly outperform those using OCR-text ($\S$4.5). Additionally, we evaluate the retrieval performance of 6 text, 4 visual, and 4 hybrid retrievers, in both pure retrieval ($\S$4.7) and end-to-end RAG ($\S$4.8) mode. The results further highlight the challenges of extracting relevant multimodal quotes from long documents. In summary, our contributions are:
@@
  • We propose MMDocRAG benchmark ($\S$2) for evaluating multimodal generation on DocVQA/RAG tasks. Our dataset include over 4,000 QA pairs, diverse forms of evidence, a mixture of gold and noisy quotes to enable nuanced quote selection, and answers with interleaved multimodal content.
@@
  • We conduct extensive evaluations ($\S$4) on multimodal RAG, covering (i) retrieval performance on 6 text, 4 visual,4 hybrid retrievers, (ii) quote selection F1 and (iii) multimodal answer quality across 37 open-source and 23 proprietary models, and 9 models finetuned using MMDocRAG dev-set.
@@ 2.1 Construction
We utilize the document corpus from the MMDocIR dataset [19], which consists of 313 documents spanning over 10 diverse domains. These documents are sufficiently long (averaging 65.1 pages) and provide rich multimodal coverage. We process the documents with MinerU [77], which leverages LayoutLMv3 [32] to detect page layouts and classify them as body text, titles, equations, figures, tables, etc. Each identified layout serves as a content-aware chunk [17], or “quote”. Text quotes correspond to layouts such as equations or paragraphs, and are stored in text format. For image quotes (e.g., tables or figures), we extract text using OCR [71] (“OCR-text”) and generate detailed descriptions using VLMs [55, 60] (“VLM-text”). Consequently, each image quote is stored in three formats: original image, OCR-text, and VLM-text. After indexing all documents, we carefully select pages with rich multimodal and text information. This process yields 2,373 high-quality pages, forming the basis for subsequent annotation.
@@ 3 Task Definition
Let $\mathcal{D}$ denote a document corpus consisting of text quotes $\mathcal{T}=\{t_{1},t_{2},\dots,t_{m}\}$ and image quotes $\mathcal{I}=\{i_{1},i_{2},\dots,i_{n}\}$, as extracted via layout detection (see Section 2.1). On average, documents in MMDocRAG contain 63 image quotes and 536 text quotes. The objective is to retrieve a subset of quotes that are most relevant to a query $Q$ from $\mathcal{T}$ and $\mathcal{I}$, by ranking them based on similarity scores, $\textrm{Sim}(Q,t)$ and $\textrm{Sim}(Q,i)$. The top-$k$ quotes, where $k\ll n+m$, are selected as candidate evidence.
@@ 4.1 Evaluation Metric
4.2 Baseline Models
Quotes Retrieval.
Multimodal Answer Generation.
We evaluate 60 latest models by using quotes as: (i) multimodal inputs for VLM, and (ii) pure-text inputs for VLM and LLM (see Appendix C.1 for implementation details). Then, we evaluate 9 finetuned models (Qwen2.5 LLMs [65] with 3, 7, 14, 32, and 72B parameters, Qwen2.5-VL VLMs [3] of 3B and 7B parameters, and InternVL-3 VLMs [90] of 8B and 9B parameters) using MMDocRAG dev-set. See Appendix C.2 for finetuning details).
4.3 Main Results
[…]
Method Metric Tokens Quote Selection Multimodal Answer Quality
Prec Rec F1 Prec Rec F1 geL ency Qlty. Coher. Logic uality
Qwen2.5-VL-3B-Inst 8.7k 265 42.5 0.5 1.0 22.8 3.0 5.3 1.2 0.105 0.283 4.07 1.07 1.49 2.45 2.17 2.25
- After Fine-tuning 8.7k 243 74.1 64.0 68.6 52.2 5.4 9.8 55.5 0.186 0.341 4.40 3.03 3.44 3.15 2.75 3.36
Qwen2.5-VL-7B-Inst 8.7k 128 58.0 14.5 23.2 31.3 11.0 16.3 16.6 0.069 0.273 4.05 1.75 1.89 2.36 2.29 2.47
- After Fine-tuning 8.7k 249 76.6 68.7 72.4 44.0 3.3 6.5 58.6 0.199 0.355 4.57 3.26 3.70 3.48 3.19 3.64
- After Fine-tuning 17.1k 268 75.4 68.3 71.7 57.3 8.2 14.4 58.8 0.205 0.356 4.05 3.68 3.75 3.51 3.58 3.71
- After Fine-tuning 17.2k 283 77.7 68.7 72.9 60.6 10.8 18.4 60.3 0.210 0.362 4.22 3.81 3.91 3.72 3.70 3.87
Figure 5: Performance difference: base/finetuned models.
Method In-token Usage Quote Sel. F1 Answer Avg.
Multimodal:$M$ Pure-Text:PT $M$ PT $\Delta\%$ $M$ PT $\Delta\%$ $M$ PT $\Delta\%$
Use the same VLM to process both multimodal and pure-text inputs.
Gemini-1.5-Pro 3.8k 3.6k -5.3 59.3 56.2 -5.2 2.77 3.03 +9.4
Gemini-2.0-Pro 3.8k 3.6k -5.3 62.0 62.8 +1.3 3.31 3.51 +6.0
Gemini-2.0-Flash 3.8k 3.6k -5.3 60.0 54.4 -9.3 3.11 3.19 +2.6
Gemini-2.0-Flash-Think 3.8k 3.6k -5.3 66.2 61.0 -7.9 3.61 3.47 -3.9
Gemini-2.5-Pro 3.7k 3.6k -2.7 65.4 65.1 -0.5 3.88 3.79 -2.3
Gemini-2.5-Flash 3.7k 3.6k -2.7 62.4 59.5 -4.6 3.76 3.55 -5.6
Claude-3.5-Sonnet 7.8k 3.8k -51.3 62.5 57.4 -8.2 3.65 3.60 -1.4
GPT-4o-mini 8.5k 3.4k -60.0 56.3 56.6 +0.5 3.60 3.70 +2.8
GPT-4o 6.4k 3.4k -46.9 62.6 57.2 -8.6 3.74 3.69 -1.3
GPT-4.1-nano 14.2k 3.4k -76.1 29.0 40.8 +40.7 3.12 3.39 +8.7
GPT-4.1-mini 9.8k 3.4k -65.3 61.2 61.0 -0.3 3.92 3.90 -0.5
GPT-4.1 6.6k 3.4k -48.5 70.2 68.3 -2.7 4.14 4.07 -1.7
InternVL3-8B 17.1k 3.6k -78.9 37.0 48.1 +30.0 3.14 3.19 +1.6
InternVL3-9B 17.2k 4.0k -76.7 50.9 45.4 -10.8 3.12 3.30 +5.8
InternVL3-14B 17.1k 3.6k -78.9 49.9 49.9 +0.0 3.29 3.48 +5.8
InternVL3-38B 17.1k 3.6k -78.9 53.9 55.0 +2.0 3.45 3.61 +4.6
InternVL3-78B 17.1k 3.6k -78.9 59.8 56.4 -5.7 3.42 3.56 +4.1
Llama4-Scout-17Bx16E 11.6k 3.3k -71.6 38.9 48.2 +23.9 3.13 3.12 -0.3
Llama4-Mave-17Bx128E 11.6k 3.3k -71.6 58.6 58.3 -0.5 3.67 3.59 -2.2
Use separate VLM/LLM to process multimodal/pure-text inputs, respectively.
Qw-VL-Plus Qw-Plus 7.1k 3.6k -49.3 25.2 55.4 +120 2.36 3.63 +53.8
Qw-VL-Max Qw-Max 7.1k 3.6k -49.3 46.8 58.9 +25.9 3.35 3.77 +12.5
QVQ-Max QwQ-Plus 6.8k 3.6k -47.1 12.3 59.6 +385 3.36 3.63 +8.0
Qw2.5-VL-7B Qw2.5-7B 7.1k 3.6k -49.3 16.6 45.8 +176 2.47 3.34 +35.2
Qw2.5-VL-32B Qw2.5-32B 7.0k 3.6k -48.6 36.2 58.9 +62.7 3.73 3.63 -2.7
Qw2.5-VL-72B Qw2.5-72B 7.1k 3.6k -49.3 57.5 59.1 +2.8 3.47 3.75 +8.1
Table 4: Using 20 quotes for multimodal generation. $\Delta\%$ is calculated by values (PT-MM)/$M$ in percentage.
4.4 Multimodal vs Pure-text Quotes: Comparison and Analysis
As shown in Table 4 and Table 9, we compare model performance when quotes are provided as either pure-text or multimodal inputs. Multimodal quotes significantly increase token usage, as images are typically encoded with more tokens. Interestingly, Gemini models maintain similar token usage across both modes, indicating efficient image encoding via visual token compression. Gemini, Claude, and GPT models demonstrate superior quote selection performance in the multimodal setting and comparable answer quality across both input types. In contrast, Qwen models perform significantly better in both quote selection and answer generation when using pure-text inputs. Smaller VLMs, compared to their LLM counterparts, struggle to effectively process long multimodal input sequences. For instance, the Qwen-7B and 32B LLMs achieve 175.9% and 62.7% higher F1 scores for quote selection, respectively, compared to their equivalent VLMs. Further qualitative analysis is provided in Appendix F.2.
Method Metric Image Quote F1 Answer Avg.
VLM OCR $\Delta$ VLM OCR $\Delta$
15 Quotes Qwen2.5-7B-Inst 59.8 49.6 -10.2 3.37 3.15 -0.22
Llama3.1-8B-Inst 61.1 52.4 -8.7 3.33 3.19 -0.14
Llama3.3-70B-Inst 71.8 64.1 -7.7 3.14 3.08 -0.06
Qwen2.5-72B-Inst 73.3 65.7 -7.6 3.76 3.55 -0.21
Qwen-Max 74.4 65.7 -8.7 3.77 3.63 -0.14
Deepseek-V3 76.5 70.2 -6.3 3.75 3.68 -0.07
Gemini-2.0-Pro 77.4 74.9 -2.5 3.50 3.41 -0.09
Gemini-2.0-Fl-Tk 74.9 73.0 -1.9 3.51 3.46 -0.05
GPT-4o 71.6 69.4 -5.9 3.73 3.65 -0.08
Avg. results 71.7 65.0 -6.7 3.54 3.42 -0.12
20 Quotes Qwen2.5-7B-Inst 53.5 43.5 -10.0 3.34 3.15 -0.19
Llama3.1-8B-Inst 52.2 45.8 -6.4 3.25 3.16 -0.09
Llama3.3-70B-Inst 65.1 60.6 -4.5 3.24 3.13 -0.11
Qwen2.5-72B-Inst 68.0 59.7 -8.3 3.75 3.50 -0.25
Qwen-Max 69.3 59.9 -9.4 3.77 3.62 -0.15
Deepseek-V3 71.8 65.8 -6.0 3.74 3.59 -0.15
Gemini-2.0-Pro 77.0 70.9 -6.1 3.51 3.32 -0.18
Gemini-2.0-Fl-Tk 73.5 68.3 -5.2 3.47 3.41 -0.06
GPT-4o 66.4 63.8 -2.6 3.69 3.59 -0.10
Avg. results 66.3 59.8 -6.5 3.53 3.39 -0.14
4.5 Multimodal Quotes as text: VLM-text vs OCR-text
4.6 Quotes Selection Analysis
Figure 7: Quotes selection accuracy at all positions.
Method Recall@10 Recall@15 Recall@20
Txt Img Txt Img Txt Img
Text DPR 25.5 53.5 31.4 59.9 35.9 63.9
ColBERT 37.4 64.1 42.8 69.1 46.0 72.8
BGE 38.8 64.9 43.6 70.2 47.0 74.2
E5 41.7 63.5 46.4 69.1 49.5 73.7
Contriever 37.7 64.1 42.8 69.8 46.8 73.4
GTE 38.2 63.0 43.3 69.1 47.3 72.8
Visual DSEwiki-$s$ 24.7 67.1 29.6 75.3 33.4 79.9
DSEdocmatix 25.6 65.7 30.1 75.0 33.7 78.2
ColPali 27.3 68.2 32.6 77.5 35.2 81.2
ColQwen 28.5 70.8 33.7 79.2 36.0 84.3
Hybrid ColP+ColB 38.2 67.3 42.6 79.2 46.8 83.4
ColP+BGE 39.2 67.3 43.5 79.2 47.7 83.4
ColQ+ColB 38.2 68.5 42.6 81.0 46.8 85.2
ColQ+BGE 39.2 68.5 43.5 81.0 47.7 85.2
Table 6: Retrieval Results.
4.7 Quotes Retrieval Results
Model Retriever Query Quote Retrieval Rec. Quote Selection F1 Multimodal Answer Quality
Text Image All Text Image All Bleu RougeL LLM-Judge
GPT-4.1 perfect - - - - 52.0 80.7 70.2 0.157 0.313 4.14
GPT-4.1 BGE original 34.6 77.8 71.0 34.2 60.8 54.4 0.137 0.299 3.53
GPT-4.1 BGE clauses 42.1 83.6 78.9 37.9 64.0 57.5 0.141 0.302 3.71
GPT-4.1 multiple clauses 49.5 86.8 84.9 41.4 65.6 59.9 0.141 0.303 3.79
Gemini2.5-Flash perfect - - - - 46.1 76.2 62.4 0.139 0.284 3.76
Gemini2.5-Flash BGE original 34.6 77.8 71.0 27.5 55.6 47.7 0.124 0.280 3.21
Gemini2.5-Flash BGE clauses 42.1 83.6 78.9 30.9 59.2 50.4 0.125 0.281 3.39
Gemini2.5-Flash multiple clauses 49.5 86.8 84.9 34.3 60.3 51.8 0.124 0.281 3.42
Table 7: End-to-end RAG Results.
4.8 End-to-end Multimodal RAG Analysis
While previous experiments employ fixed quotes to isolate generation evaluation, real-world RAG systems must contend with imperfect retrieval. To bridge this gap and assess the robustness of multimodal generation under realistic conditions, we conduct end-to-end experiments that jointly evaluate retrieval and generation performance.
Experimental Setup.
We evaluate four retrieval configurations with varying degrees of retrieval quality: (1) Perfect retriever (upper bound): All gold quotes provided alongside noisy quotes, maintaining the 20-quote setting (8 images, 12 texts). Single retriever: BGE [83] with either (2) original questions or (3) expanded multi-clause queries. (4) Multi-retriever ensemble: Combination of BGE, Qwen3-0.6B [87], BM25 [68], and E5 [78] retrievers with query expansion. Note that for multi-clause queries and multi-retriever methods, we consolidate top quotes from multi-retriever/clause via reranking using Qwen3-0.6B-reranker [87]. We focus on single-vector embedding models for compatibility with production vector databases (e.g., Milvus), excluding multi-vector approaches like ColBERT [37] and ColQwen [21].
Results and Analysis.
Table 7 presents the end-to-end performance across retrieval configurations. Three key observations emerge: (i) Retrieval-generation correlation: A clear positive correlation exists between retrieval recall and downstream performance. When retrieval recall drops from perfect (100%) to 71.0% using single BGE with original queries, GPT-4.1’s quote selection F11 degrades from 70.2 to 54.4 (-22.5%), while answer quality drops from 4.14 to 3.53 (-14.7%). (ii) Query expansion benefits: Expanding queries into multi-clause formulations consistently improves retrieval recall (+7.9% absolute for BGE), which cascades into better generation performance. This suggests that comprehensive query understanding remains crucial for document-grounded multimodal generation. (iii) Multi-retriever robustness: Ensemble approaches achieve substantially higher recall (84.9%) compared to single retrievers (71.0-78.9%), narrowing the performance gap with perfect retrieval. Even leading models like GPT-4.1 and Gemini-2.5-Flash experience approximately 10% performance degradation under realistic retrieval conditions, highlighting the continued challenge of end-to-end multimodal RAG. These findings validate that MMDocRAG effectively captures the cascading challenges in practical RAG systems, where imperfect retrieval directly impacts the quality of multimodal answer generation. The benchmark thus serves as a comprehensive testbed for both retrieval and generation components in document-grounded multimodal systems.
5 Related Work
Interleaved Text-Image Generation
Multimodal RAG and Benchmarks.
Retrieval-Augmented Generation (RAG) retrieves relevant quotations as context for answer generation [39, 86]. Multimodal RAG (MRAG) extends RAG by retrieving and leveraging multimodal knowledge (e.g., image-text pairs) for VQA [5, 43]. MuRAR [91] tackles source attribution by retrieving multimodal elements from webpage. M2RAG [48] builds upon MuRAR by proposing a multi-stage image insertion framework that uses model multiple times during answer generation. Although MuRAR and M2RAG enable multimodal answer generation, their benchmarks are limited to webpage domain and lack annotations of supporting evidence.
DocVQA and DocRAG Benchmarks.
Early DocVQA benchmarks focus on single-page VQA, such as DocVQA [49], InfoVQA [50], and TAT-DQA [89]. To mitigate the limitation of single-page input, DUDE [38], MP-DocVQA [76], SildeVQA [72] extend context lengths to averages 5.7, 8.3, and 20 pages respectively. However, these tasks are yet to be document-level [16] reasoning or understanding. Two most recent MMLongBench-Doc [47] and DocBench [92], formulate DocVQA as long-context tasks by inputting entire documents (averaging 50-70 pages). To address increasing document length, M3DocVQA [10], M-Longdoc [9], and MMDocIR [19] propose DocRAG tasks, incorporating evidence retrieval followed by answer generation over the retrieved multimodal evidence. To the best of our knowledge, no existing DocVQA or DocRAG benchmarks focus on multimodal interleaved generation.
6 Conclusion
Appendix Overview
The appendix includes the following sections:
  • Appendix B: Provides supplementary experimental results, with (B.1) related results by using 15 quotes for multimodal generation, (B.2) comprehensive results comparing thinking and non-thinking modes, (B.3) comprehensive results comparing different models using OCR and LLM text for image quote representation, and fine-grained results by document type (B.4), question type (B.5), and evidence type (B.6).
Appendix A Evaluation Metric of Multimodal Answer Generation
A.1 Related Work of multimodal generation
Multimodal generation, particularly interleaved image-text sequence generation, involves generating outputs that integrate visual and textual information in a cohesive manner (see Section 5). This capability facilitate applications such as storytelling, question answering, and document comprehension. Recent benchmark, MM-Interleaved [75], MMIE [81], GATE Opening [88], and M2RAG [48] provide comprehensive evaluations for multimodal generation. Commonly adopted metrics include fluency, relevance, image-text coherence, and content quality. These are evaluated through human annotation or automated scoring using large language models such as GPT-4. Specifically, fluency assesses the grammatical correctness and readability of text, relevance measures the alignment of generated content with the prompt, image-text coherence evaluates the logical connection between images and text, and content quality addresses the completeness and richness of the output. Our benchmark, MMDocRAG, adopts established metrics such as fluency, image-text coherence, and content quality. Additionally, we incorporate BLEU [59] and ROUGE-L [42] scores to quantitatively assess the lexical similarity between generated and gold answers.
A.2 Evaluation Metrics: details and implementations
1. Quote Selection Metrics.
2. Surface-level Similarity Metrics.
BLEU $=BP\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right),$ $\text{where }BP$ $=\begin{cases}1,&\text{if }c>r\\ \exp\left(1-\frac{r}{c}\right),&\text{if }c\leq r\end{cases}$ (2)
$\text{ROUGE-L}=\frac{(1+\beta^{2})\cdot R_{LCS}\cdot P_{LCS}}{R_{LCS}+\beta^{2}P_{LCS}},\quad P_{LCS}=\frac{LCS(\text{Gen},\text{Ref})}{|\text{Gen}|},\ R_{LCS}=\frac{LCS(\text{Gen},\text{Ref})}{|\text{Ref}|}$ (3)
where $R_{LCS}$ and $P_{LCS}$ are the recall and precision based on LCS length, and $|\cdot|$ refers to the length of generated or reference answer. $\beta$ is typically set to favor recall ($\beta=1.2$ by default).
3. LLM-as-Judge Evaluation Criteria.